# On Efficient Variants of Segment Anything Model: A Survey

### v1

@@ Abstract
The Segment Anything Model (SAM) is a foundational model for image segmentation tasks, known for its strong generalization across diverse applications. However, its impressive performance comes with significant computational and resource demands, making it challenging to deploy in resource-limited environments such as mobile devices. To address this, a variety of SAM variants have been proposed to enhance efficiency without sacrificing accuracy. This survey provides the first comprehensive review of these efficient SAM variants. We begin by exploring the motivations driving this research. We then present core techniques used in SAM and model acceleration. This is followed by an in-depth analysis of various acceleration strategies, categorized by approach. Finally, we offer a unified and extensive evaluation of these methods, assessing their efficiency and accuracy on representative benchmarks, and providing a clear comparison of their overall performance.
@@ 1 Introduction
The emergence of foundation models [8] has led to a thorough revolution in the field of artificial intelligence (AI). Foundation models are large-scale neural networks pre-trained on massive data, which have powerful representing ability and strong generalization to perform on various tasks. In the field of natural language processing (NLP), the recently popular research trend is about the large language models (LLMs) [160, 116], which has been witnessed great development with prominent works like OpenAI’s GPT series [9, 1], Google’s PaLM series [21, 3] and Meta’s LLaMA series [111, 112]. Meanwhile, the success of Vision Transformer (ViT) [29] which firstly introduces the Transformer [115] architecture into the field of computer vision (CV), has brought up a new era for vision foundation models (VFMs). Vision-Language foundation models like CLIP [89], LLaVA [67], and Video-ChatGPT [79] etc., which aim at aligning the vision and language modalities, have demonstrated promising performance on plenty of downstream vision tasks[58, 34, 167, 110].
@@
Despite its success in a wide range of applications, the original Segment Anything Model (SAM), particularly SAM-H, faces significant limitations due to its slow runtime and high computational cost. These challenges become even more pronounced when SAM is deployed in resource-constrained or real-time environments, such as edge devices and mobile applications. As the demand for deploying machine learning models in practical, resource-constrained scenarios continues to grow, SAM’s current design proves inefficient for widespread use. This has led to a pressing need for more lightweight and efficient variants that can maintain the model’s powerful segmentation capabilities while addressing these constraints. The challenge of optimizing SAM for efficiency is further amplified by the increasing emphasis on real-time applications, mobile platforms and embedded systems, where computational resources are limited. As the community strives to overcome these obstacles [161, 114, 147, 103, 118, 164, 77, 18], a comprehensive understanding of the latest advancements in making SAM more efficient is crucial. Consequently, it is both timely and necessary to conduct a detailed survey of the ongoing efforts aimed at improving SAM’s efficiency and extending its applicability across diverse environments.
With the growing research related to SAM, several surveys have been presented to provide an overview from different perspectives [150, 149, 152, 157, 151, 155]. However, these existing surveys primarily focus on SAM’s downstream applications and exhibit several limitations: 1) None of them address the emerging field of improving SAM’s efficiency, which is gaining significant attention and is crucial for practical deployment in real-world applications. 2) With the exception of [151], these surveys lack a structured taxonomy that would allow for clearer organization and reference. 3) Most prior surveys focus on collecting and describing SAM-based methods but lack a systematic evaluation or comparison of these approaches. To address these gaps, we conduct this survey, not only to comprehensively review the development of efficient Segment Anything Models but also to fairly evaluate and compare them. The main contributions of this work can be summarized as follows:
@@
  • We offer comprehensive evaluations and comparisons of the efficiency and accuracy of these variants, aiming to assist researchers in selecting models that best meet their performance and application needs.
@@
The rest of this survey is organized as follows. In Section 2, we begin by introducing the background of the original SAM, followed by a review of efficient backbones for visual representation and model compression techniques that can be applied to enhance SAM’s efficiency. In Section 3, we categorize the existing methodologies based on their objectives and techniques, providing a detailed review of each category. We also discuss several potential research directions for further accelerating SAM. In Section 4, we conduct a fair evaluation of these models in terms of efficiency, accuracy, and the resulting trade-offs. Finally, in Section 5, we provide a brief summary of this survey.
@@ 2 Preliminaries
In this section, we will first introduce the details and applications of the Segment Anything Model in Section 2.1. In Section 2.2, we review efficient backbones that could potentially replace SAM’s image encoder, and in Section 2.3, we examine model compression techniques that could be applied to SAM. For a more comprehensive understanding of efficient backbones and model compression methods, we recommend referring to surveys [86, 132].
@@ 2.1.1 Model
SAM consists of three components: an image encoder, a prompt decoder and a mask decoder, as shown in Fig. 1 (a). The image encoder is a MAE [44] pre-trained Vision Transformer(ViT) with minimal modification. It takes the preprocessed images as input and outputs an image embedding for each image. The prompt decoder is to embed prompts: points, boxes, masks and text. The two embeddings are then fed into the light-weight mask decoder, which is built upon two modified Transformer decoder blocks [115] and a few prediction heads, to generate valid masks. Based on SAM’s architecture, SAM 2 additionally introduces a streaming memory mechanism. Specifically, a memory encoder, a memory bank and a memory attention module. The structure of SAM 2 is illustrated in Fig. 1 (b). With the memory mechanism, SAM 2 can handling videos frame by frame. The memory encoder is to generate memory of the prediction of the current frame and send it to the memory bank. The memory bank stores memories of recent prediction, feature maps of prompted frames and high-level semantic information of the target objects (i.e. object pointers). The memory attention mechanism is to make the image embedding from image encoder fully interact with the information from memory bank, resulting in the refined embedding. Apart from the memory mechanism, SAM 2 also adopts the MAE pre-trained Hiera [96] as image encoder which is more efficient than ViT-H, with expectation for faster speed.
@@ 2.1.3 Application
One of the major applications of SAM is medical image segmentation. According to [157], works in this area can be classified into two types. Some aim at testing SAM’s segmentation performance in CT images [95, 46], MRI images [82], pathological images [25], and so on. Others focus on improving the adaptation of SAM to these tasks by fine-tuning [78, 64], auto-prompting [48, 98] or framework modification [156]. Furthermore, works like [57, 35] attempt to improve medical SAM methods’ efficiency. SAM is also applied to object detection across a diverse range of real-world scenes [149, 51], including crack detection [2] and crater detection [36] in civil infrastructure defect assessment, crop disease and pest detection [63] in agriculture, anomaly detection [11] and remote sensing [85, 119]. Moreover, Segment Anything has been adapted to Edit Everything [130], Inpaint Anything [141] and Caption Anything [121] to handle image editing tasks.
Apart from image segmentation tasks, SAM has been widely applied to various video tasks [151]. A large portion of the works concentrate on two basic tasks: video object segmentation (VOS) [154, 158, 19] and video object tracking (VOT) [137, 20, 92]. Researchers also explore SAM’s application in tasks related to generation, for example, video super-resolution [76] and video dataset annotation generation [42, 24]. Besides these, SAM is further utilized as an intermediate tool in video editing tasks [163, 126]. Beyond 2D tasks, SAM is also extended to 3D vision filed. [100] applies SAM to 3D reconstruction and [28] applies it to 3D point cloud segmentation. The recent work [134] aims at achieving real-time segmentation for any 3D thing in an online setting. For the recent proposed SAM 2, there are already some studies exploring its application in both image and video tasks. A popular trend is to apply SAM 2 to medical images and video tasks [155]. Works like [99, 27, 135] evaluate SAM 2’s performance in medical images in both 2D and 3D modalities, while some others [140, 75] test it on surgical video segmentation tasks. Researchers are also seeking for strategies to adapt SAM 2 to medical tasks better [136, 166]. Besides these, SAM 2 has been applied to some specific image segmentation tasks like digital pathology semantic segmentation [153], mesh part segmentation [109] and solar panels segmentation [91]. Moreover, several works [113, 70] have utilized SAM 2 on the challenging Large-scale Video Object Segmentation (LSVOS) task and achieved good results.
@@ 2.1.4 Limitation
Although SAM has demonstrated promising performance across various tasks, it still faces two key challenges in practical applications: 1) it often struggles to predict complete masks for fine structures, leading to imprecise boundaries; and 2) it is not real-time and remains resource-intensive, particularly when using heavy image encoders like ViT-H. To address these issues, works such as [54, 49] aim to improve mask quality by utilizing high-resolution images, while others [161, 147, 164, 131] focus on creating more efficient architectures to reduce SAM’s time and resource consumption. Previous surveys [78, 158, 100] have explored recent advances in enhancing SAM for higher-quality results. In this survey, we focus specifically on efforts to improve SAM’s efficiency.
@@ 2.2.1 Efficient Vision Transformer
To reduce computational cost from a structural perspective, MobileViT [80], a pioneering hybrid architecture, creatively integrates CNN blocks (MobileNetV2 blocks [97]) with Transformer blocks into a single model. Subsequent works such as [127, 62, 10] basically follow this idea to build up efficient ViTs with hybrid structures, which have been widely used to substitute SAM’s heavy image encoder. In [161, 103], TinyViT [127] serves as the efficient backbone, while in [162] and [159], EfficientFormerV2 [62] and EfficientViT [10] supplant SAM’s original image encoder, respectively. Another influential ViT design, MetaFormer [142], which abstracts the attention mechanism into a broader concept called the token mixer, can deliver performance on par with Transformers using various token mixers. The simplest variant, PoolFormer, which uses pooling operations as the token mixer without introducing additional learnable parameters, has been leveraged as the base architecture to develop the light-weight image encoder for Lite-SAM [33].
Researchers have also made significant progress in optimizing the attention mechanism. It has been observed that the softmax operation within the attention mechanism contributes significantly to the overall computational cost. In EfficientViT [10], a novel ReLU linear attention is proposed to achieve global receptive field with higher efficiency. This efficient backbone is further adopted in [159] to accelerate SAM. Improvements to the attention mechanism have also been made at the hardware level. FlashAttention [23] significantly reduces computation costs through techniques such as tiling, kernel fusion, and recomputation. And it is utilized in SAM acceleration works [101, 104] to reduce memory need and enhance computation efficiency.
@@ 2.2.2 Transformer-alternative Models
where $r,k,v$ represents the shifted tokens of receptance, key and value respectively and $W$ refers to weight.
@@ 2.3.2 Quantization
Quantization is the process to convert model’s high precision weights/activation values $X$ (e.g. 32-bit floating point) to low precision formats $X_{q}$ (e.g. 16-bit floating point, 8-bit integer). One of the widely used quantization function, uniform symmetric quantization, is formulated as follows,
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=16em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, where level=4text width=18em,font=,, where level=5text width=12em,font=,, [Efficient SAM variants, ver [Accelerating SegAny3.1), fill=green!10 [Training from Scratch3.1.1), fill=blue!10 [SAM-different architectures, fill=yellow!10 [FastSAM [161], fill=gray!10] [SqueezeSAM [114], fill=gray!10] ] [SAM-like architectures, fill=yellow!10 [EfficientSAM [131], fill=gray!10] [RAP-SAM [133], fill=gray!10] [SAM 2 [93], fill=gray!10] ] ] [Knowledge Distillation3.1.2), fill=blue!10 [With Light-weight ViT Encoders, fill=yellow!10 [MobileSAM [147], fill=gray!10] [ESAM [162], fill=gray!10] [TinySAM [103], fill=gray!10] ] [With Pure CNN Encoders, fill=yellow!10 [NanoSAM [84], fill=gray!10] [RepViT-SAM [118], fill=gray!10] [EdgeSAM [164], fill=gray!10] ] [With Attention-modified Encoders, fill=yellow!10 [EfficientViT-SAM [159], fill=gray!10] [FastSAM3D [101], fill=gray!10] [SAM-Lightening [104], fill=gray!10] [RWKV-SAM [87], fill=gray!10] ] ] [Quantization3.1.3), fill=blue!10 [TinySAM [103], fill=gray!10] [PTQ4SAM [77], fill=gray!10] ] [Pruning3.1.4), fill=blue!10 [SlimSAM [18], fill=gray!10] ] [Code Refactorization3.1.5), fill=blue!10 [SAMFast [88], fill=gray!10] ] ] [Accelerating SegEvery3.2), fill=green!10 [,coordinate, [FastSAM [161], fill=gray!10, tier=number] ] [Efficient sampling strategy, fill=blue!10 [MobileSAMV2 [148], fill=gray!10, tier=number] [TinySAM [103], fill=gray!10] [LiteSAM [33], fill=gray!10] ] ] ]

### v2

@@ Abstract
The Segment Anything Model (SAM) is a foundational model for image segmentation tasks, known for its strong generalization across diverse applications. However, its impressive performance comes with significant computational and resource demands, making it challenging to deploy in resource-limited environments such as edge devices. To address this, a variety of SAM variants have been proposed to enhance efficiency while keeping accuracy. This survey provides the first comprehensive review of these efficient SAM variants. We begin by exploring the motivations driving this research. We then present core techniques used in SAM and model acceleration. This is followed by a detailed exploration of SAM acceleration strategies, categorized by approach, and a discussion of several future research directions. Finally, we offer a unified and extensive evaluation of these methods across various hardware, assessing their efficiency and accuracy on representative benchmarks, and providing a clear comparison of their overall performance.
@@ 1 Introduction
The emergence of foundation models [8] has led to a thorough revolution in the field of artificial intelligence. Foundation models are large-scale neural networks pre-trained on massive data, which have powerful representing ability and strong generalization to perform various tasks. In the field of natural language processing, the recently popular research trend is about the large language models (LLMs) [172, 128], which has witnessed great development with prominent works like OpenAI’s GPT series [9, 1], Google’s PaLM series [23, 3] and Meta’s LLaMA series [123, 124]. Meanwhile, the success of Vision Transformer (ViT) [31] which firstly introduces the Transformer [127] architecture into the field of Computer Vision, has brought up a new era for vision foundation models (VFMs). Vision-language foundation models like CLIP [101], LLaVA [76], and Video-ChatGPT [89] etc., which aim at aligning the vision and language modalities, have demonstrated promising performance on plenty of downstream vision tasks[65, 37, 180, 122].
@@
Despite its success in a wide range of applications, the original Segment Anything Model, particularly SAM-H, faces significant limitations due to its slow runtime and high computational cost. These challenges become even more pronounced when SAM is deployed in resource-constrained or real-time environments. As the demand for deploying machine learning models in practical, resource-constrained scenarios continues to grow, SAM’s current design proves inefficient for widespread use. This has led to a pressing need for more lightweight and efficient variants that can maintain the model’s powerful segmentation capabilities while addressing these constraints. As the community strives to overcome these obstacles [173, 126, 159, 115, 130, 176, 87, 20], a comprehensive understanding of the latest advancements in making SAM more efficient is crucial. Consequently, it is both timely and necessary to conduct a detailed survey of the ongoing efforts aiming at improving SAM’s efficiency and extending its applicability across diverse environments.
With the growing research related to SAM, several surveys have been presented to provide an overview from different perspectives [162, 161, 164, 169, 163, 167]. However, these existing surveys primarily focus on SAM’s downstream applications and exhibit several limitations: 1) None of them address the emerging field of improving SAM’s efficiency, which is gaining significant attention and is crucial for practical deployment in real-world applications. 2) With the exception of [163], most surveys lack a structured taxonomy that would allow for clearer organization and reference. 3) Most prior surveys focus on collecting and describing SAM-based methods but lack a systematic evaluation or comparison of these approaches. To address these gaps, we conduct this survey, not only to comprehensively review the development of efficient Segment Anything Models, but also to fairly evaluate and compare them. The main contributions of this work can be summarized as follows:
@@
  • We offer comprehensive evaluations and comparisons of the efficiency and accuracy of these variants across different hardware, aiming at assisting researchers in selecting models that best meet their performance and application needs.
@@
The rest of this survey is organized as follows. In Section 2, we begin by introducing the background of the original SAM, followed by a review of model acceleration techniques that can be applied to enhance SAM’s efficiency. In Section 3, we categorize the existing methodologies based on their objectives and techniques, providing a detailed review of each category. We also discuss several potential research directions for further accelerating SAM. In Section 4, we conduct unified evaluations of these models in terms of efficiency and accuracy. Finally, in Section 5, we provide a summary of this survey.
@@ 2 Preliminaries
In this section, we will first introduce the details and applications of the Segment Anything Model (SAM) in Section 2.1. In Section 2.2, we briefly review efficient backbones that could potentially replace SAM’s image encoder, and in Section 2.3, we discuss model compression techniques that could be applied to SAM. For a more comprehensive understanding of efficient backbones and model compression methods, we recommend referring to surveys [98, 144].
@@ 2.1.1 Model
SAM consists of three components: an image encoder, a prompt decoder, and a mask decoder, as shown in Fig. 1 (a). The image encoder is an MAE [50] pre-trained Vision Transformer(ViT) with minimal modification. It takes the preprocessed images as input and outputs an image embedding for each image. The prompt decoder is to embed prompts: points, boxes, masks, and text. The two embeddings are then fed into the lightweight mask decoder, which is built upon two modified Transformer decoder blocks [127] and a few prediction heads, to generate valid masks. Based on SAM’s architecture, SAM 2 additionally introduces a streaming memory mechanism. Specifically, a memory encoder, a memory bank, and a memory attention module. The structure of SAM 2 is illustrated in Fig. 1 (b). With the memory mechanism, SAM 2 can handle videos frame by frame. The memory encoder generates the memory of the current frame and sends it to the memory bank. The memory bank stores memories of recent predictions, feature maps of prompted frames, and high-level semantic information of the target objects (i.e. object pointers). The memory attention mechanism aims to make the image embedding from the image encoder fully interact with the information from the memory bank, resulting in refined embedding. Apart from the memory mechanism, SAM 2 also adopts the MAE pre-trained Hiera [108] as the image encoder which is more efficient than ViT-H, with expectation for faster speed.
@@ 2.1.3 Application
One of the major applications of SAM is medical image segmentation. According to [169], works in this area can be classified into two types. Some aim at testing SAM’s segmentation performance in CT images [107, 53], MRI images [93], pathological images [27], and so on. Others focus on improving the adaptation of SAM to these tasks by fine-tuning [88, 71], auto-prompting [55, 110] or framework modification [168]. Furthermore, works like [64, 39, 90] attempt to improve medical SAM methods’ efficiency. SAM is also applied to object detection across a diverse range of real-world scenarios [161, 58], including crack detection [2] and crater detection [40] in civil infrastructure defect assessment, crop disease and pest detection [70] in agriculture, anomaly detection [11] and remote sensing [96, 131]. Moreover, Segment Anything has been adapted to Edit Everything [142], Inpaint Anything [153], and Caption Anything [133] to handle image editing tasks.
Apart from image segmentation tasks, SAM has been widely applied to various video tasks [163]. A large portion of the work concentrates on two basic tasks: video object segmentation (VOS) [166, 170, 21] and video object tracking (VOT) [149, 22, 104]. Researchers also explore SAM’s application in tasks related to generation, for example, video super-resolution [86] and video dataset annotation generation [48, 26]. Besides these, SAM is further utilized as an intermediate tool in video editing tasks [175, 138]. Beyond 2D tasks, SAM is also extended to the 3D vision field. Shen et al. [112] applies SAM to 3D reconstruction and Dong et al. [30] applies it to 3D point cloud segmentation. The recent work [146] aims at achieving real-time segmentation for any 3D thing in an online setting. For the recently proposed SAM 2, some studies have already explored its application in both image and video tasks. A popular trend is to apply SAM 2 to medical images and video tasks [167]. Works like [111, 29, 147] evaluate SAM 2’s performance in medical images in both 2D and 3D modalities, while some others [152, 85] test it on surgical video segmentation tasks. Researchers are also seeking strategies to adapt SAM 2 to medical tasks better [148, 178]. Besides these, SAM 2 has been applied to some specific image segmentation tasks like digital pathology semantic segmentation [165], mesh part segmentation [121], and solar panels segmentation [103]. Moreover, several works [125, 80] have utilized SAM 2 on the challenging Large-scale Video Object Segmentation (LSVOS) task and achieved good results.
@@ 2.1.4 Limitation
Although SAM has demonstrated promising performance across various tasks, it still faces two key challenges in practical applications: 1) it often struggles to predict complete masks for fine structures, leading to imprecise boundaries; and 2) it is not real-time and remains resource-intensive, particularly when using heavy image encoders like ViT-H. To address these issues, works such as [61, 56] aims to improve mask quality by utilizing high-resolution images, while others [173, 159, 176, 143] focus on creating more efficient architectures to reduce SAM’s time and resource consumption. Previous surveys [88, 170, 112] have explored recent advances in enhancing SAM for higher-quality results. In this survey, we focus specifically on efforts to improve SAM’s efficiency.
@@ 2.2.1 Efficient Vision Transformer
To reduce computational cost from a structural perspective, MobileViT [91], a pioneering hybrid architecture, creatively integrates CNN blocks (MobileNetV2 blocks [109]) with Transformer blocks into a single model. Subsequent works such as [139, 69, 10] basically follow this idea to build up efficient ViTs with hybrid structures, which have been widely used to substitute SAM’s heavy image encoder. In [173, 115], TinyViT [139] serves as the efficient backbone, while in [174] and [171], EfficientFormerV2 [69] and EfficientViT [10] supplant SAM’s original image encoder, respectively. Another influential ViT design, MetaFormer [154], which abstracts the attention mechanism into a broader concept called the token mixer, can deliver performance on par with Transformers using various token mixers. The simplest variant, PoolFormer, which uses pooling operations as the token mixer without introducing additional learnable parameters, has been leveraged as the base architecture to develop the light-weight image encoder for Lite-SAM [36]. The work [73] expedites Vision Transformer by utilizing the creatively proposed token clustering and token reconstruction mechanism, which has already been applied to SAM’s image encoder, resulting in the variant Expedit-SAM [72] with faster inference speed in image segmentation tasks.
Researchers have also made significant progress in optimizing the attention mechanism. It has been observed that the softmax operation within the attention mechanism contributes significantly to the overall computational cost. In EfficientViT [10], a novel ReLU linear attention is proposed to achieve a global receptive field with higher efficiency. This efficient backbone is further adopted in [171] to accelerate SAM. Improvements to the attention mechanism have also been made at the hardware level. FlashAttention [25] significantly reduces computation costs through techniques such as tiling, kernel fusion, and recomputation. And it is utilized in SAM acceleration works [113, 116] to reduce memory need and enhance computation efficiency.
@@ 2.2.2 Transformer-alternative Models
State Space Models (SSMs), with near-linear computational complexity relative to sequence length, have emerged as promising alternatives to Transformers. Various methods [44, 46, 35] have been developed to enhance SSMs. One recent advancement, Mamba [43], has made significant strides in processing long sequences. Mamba’s core innovation is the Selective State Spaces, which allow the model to dynamically decide whether to propagate or forget information, enabling more efficient sequence modeling. Additionally, the researchers incorporated a hard-aware algorithm to further accelerate computation on GPUs. Building on this, Vision Mamba [179] applies Mamba to visual representation tasks. Vim introduces Bi-directional SSMs to better capture context within image patch sequences. With its strong performance and efficiency, Vision Mamba is emerging as a potential backbone for vision foundation models.
@@
where $r,k,v$ represents the shifted tokens of receptance, key, and value respectively, and $W$ refers to weight.
@@ 2.3.2 Quantization
Quantization is the process to convert a model’s high precision weights/activation values $X$ (e.g. 32-bit floating point) to low precision formats $X_{q}$ (e.g. 16-bit floating point, 8-bit integer). One of the widely used quantization functions, uniform symmetric quantization, is formulated as follows,
@@ 2.3.3 Pruning
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=16em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, where level=4text width=18em,font=,, where level=5text width=12em,font=,, [Efficient SAM variants, ver [Accelerating SegAny3.1), fill=green!10 [Training from Scratch3.1.1), fill=blue!10 [SAM-different architectures, fill=yellow!10 [FastSAM [173], fill=gray!10] [SqueezeSAM [126], fill=gray!10] ] [SAM-like architectures, fill=yellow!10 [EfficientSAM [143], fill=gray!10] [RAP-SAM [145], fill=gray!10] [SAM 2 [105], fill=gray!10] ] ] [Knowledge Distillation3.1.2), fill=blue!10 [With Light-weight ViT Encoders, fill=yellow!10 [MobileSAM [159], fill=gray!10] [ESAM [174], fill=gray!10] [TinySAM [115], fill=gray!10] ] [With Pure CNN Encoders, fill=yellow!10 [NanoSAM [95], fill=gray!10] [RepViT-SAM [130], fill=gray!10] [EdgeSAM [176], fill=gray!10] ] [With Attention-modified Encoders, fill=yellow!10 [EfficientViT-SAM [171], fill=gray!10] [FastSAM3D [113], fill=gray!10] [SAM-Lightening [116], fill=gray!10] [RWKV-SAM [99], fill=gray!10] ] ] [Quantization3.1.3), fill=blue!10 [TinySAM [115], fill=gray!10] [PTQ4SAM [87], fill=gray!10] ] [Pruning3.1.4), fill=blue!10 [SlimSAM [20], fill=gray!10] ] [Code Refactorization3.1.5), fill=blue!10 [SAMFast [100], fill=gray!10] ] ] [Accelerating SegEvery3.2), fill=green!10 [Panoptic Segmentation Strategy, fill=blue!10 [FastSAM [173], fill=gray!10, tier=number] ] [Efficient Sampling Strategy, fill=blue!10 [MobileSAMV2 [160], fill=gray!10, tier=number] [TinySAM [115], fill=gray!10] [LiteSAM [36], fill=gray!10] ] ] ]